# How Interpretable are Reasoning Explanations from Prompting Large Language Models?

### v2

@@ Abstract
Prompt Engineering has garnered significant attention for enhancing the performance of large language models across a multitude of tasks. Techniques such as the Chain-of-Thought not only bolster task performance but also delineate a clear trajectory of reasoning steps, offering a tangible form of explanation for the audience. Prior works on interpretability assess the reasoning chains yielded by Chain-of-Thought solely along a singular axis, namely faithfulness. We present a comprehensive and multifaceted evaluation of interpretability, examining not only faithfulness but also robustness and utility across multiple commonsense reasoning benchmarks. Likewise, our investigation is not confined to a single prompting technique; it expansively covers a multitude of prevalent prompting techniques employed in large language models, thereby ensuring a wide-ranging and exhaustive evaluation. In addition, we introduce a simple interpretability alignment technique, termed Self-Entailment-Alignment Chain-of-thought, that yields more than 70% improvements across multiple dimensions of interpretability. Code is available at https://github.com/wj210/CoT_interpretability
@@
Yeo Wei Jie1, Ranjan Satapathy2, Goh Siow Mong, Rick2, Erik Cambria1 1Nanyang Technological University 2A*STAR, IHPC yeow0082@e.ntu.edu.sg
@@ 1 Introduction
In recent trends, Large Language Models (LLM) have shown impressive performance across a diverse array of tasks, primarily through extensive scaling of model size Brown et al. (2020). Techniques such as instruct-tuning Wei et al. (2021) applied across diverse tasks have empowered LLMs to execute inference on previously unseen tasks. One attributing factor lies with the extensive efforts put into innovating new ways of prompting the LLM to better exploit their knowledge base.
Chain-of-Thought (CoT) Wei et al. (2022) has gathered much attention due to its simple setup which allows the LLM to generate not only the task output but also the steps undertaken. In addition to its efficacy in enhancing the model’s performance, this prompting method concurrently touches on one of the important aspects of utilizing these models for decision-making: interpretability.
The assumption is that the reasoning chain preceding the answer illustrates the model’s thought process, enabling the audience to understand how the answer is derived. However, such claims though seemingly plausible should be taken lightly as they may not be faithful to the model’s reasoning process Jacovi and Goldberg (2020). In this context, plausibility refers to the extent to which an explanation resonates with and is deemed acceptable by a human audience. Faithfulness, on the other hand, is characterized by the extent to which the explanation accurately reflects the model’s decision-making process.
There has been a large number of works that seek to introduce modifications to CoT, including Self-Consistency Wang et al. (2022b) and Least-to-Most Zhou et al. (2022). We introduce a simple extension to the list of CoT variants, but purely with a focus on enhancing interpretability in the reasoning chain. The approach coined Self-Entailment-Alignment CoT (SEA-CoT) similarly utilizes a form of consistency between the set of possible outcomes, with an additional touch of alignment towards desirable explainability qualities. Moreover, we conduct an extensive investigation into the reasoning explanations by evaluating under three pivotal axes of interpretability: faithfulness, robustness, and utility on three commonsense reasoning datasets. These assessments are implemented across multiple prompting techniques including CoT and various adaptations of it.
@@ 3 Prompt Techniques
Figure 2: Overview of different prompting techniques to derive the reasoning chain, to serve as the explanation (boxed with dashed line). [Left to Right]: Cot, SC-CoT, SEA-CoT, QD and Self-Refine (SR). SC-CoT and SEC-CoT differ in the explanation selection stage, where the former selects based on maximum cumulative probability and the latter on two objectives: entailment, $E$, and overlap, $O$ with an additional forward pass. Each robot figure denotes a forward pass from the LLM, SR stops when a stopping criteria is encountered or exceeds the max number of passes. SR requires the most pass, 3 per round.
@@ 4 Proposed Approach
Most adaptations on CoT are only aimed at maximizing task performance as covered in Section 3. Our work is instead focused on enhancing the interpretability of the presented reasoning chain preceding the task output. We adapt from SC-CoT, by focusing on the $N$ sequences produced, ranked based on specified objectives. Instead of picking explanations based on heuristics such as highest cumulative probability, the reasoning is chosen based on the maximization of two objectives: entailment and overlapping score between the supported context $(x\oplus\hat{y})$ and reasoning $\hat{e}$. We posit that a credible explanation should intrinsically align with the given context it aims to elucidate Jie et al. (2024); in this scenario, it encompasses both the question being addressed and the prediction label, measured by the level of entailment. We additionally maximize the overlap between two sets of key tokens [^1] The two set of tokens are compared after removing any stopwords to minimize noise within the context, which we show in later experiments to be beneficial towards producing higher quality explanations. This simple approach can be regarded as performing a self-alignment step to pick the most suitable explanation with the $N$ sequences.
@@ 5 Interpretability Qualities
Utility: A largely understudied but important trait, utility is paramount towards maximizing the information conveyed to the audience. A useful explanation can allow the discovery of new knowledge to human users such as understanding the causal relationships or enable more efficient knowledge distillation between neural models.
@@ 5.4 Counterfactual reasoning
An alternative method to ascertain faithfulness follows by evaluating whether an explanation would change when the original question is modified in a different direction, particularly when directed towards a counterfactual scenario.  Atanasova et al. (2023) shows that an instance of unfaithfulness can be detected if the counterfactual explanation, $e^{\prime}$ does not acknowledge the modifications, $c$ in the counterfactual instance $x_{i}^{\prime}:y^{\prime}$, yet still successfully predicting the counterfactual label, $y^{\prime}\neq y$. The distinction from Section 5.2 is that besides detecting signs of unfaithfulness, it also embodies a directed approach that assesses a model’s capacity to contemplate alternative scenarios. Conversely, introducing mistakes can be seen as an undirected measure aimed at gauging the decline in confidence, given an erroneous prior belief. We deemed an instance of unfaithfulness under the following conditions:
@@ 6 Experiments
Datasets: We implement the perturbation experiments across three commonsense reasoning benchmarks.
@@ 6.1 Results
The key distinction between SC-CoT and SEA-CoT is the latter’s self-critique step, which evaluates how its explanations align with the context and the intended answer. This approach significantly boosts utility and reduces unfaithfulness in counterfactual contexts. Higher utility scores support the hypothesis that context-aligned stimuli enhance the efficiency of learning signals, facilitating easier training for student models. Looking closer in Figure 6, where the word "shunned" is mentioned while other baselines used "would not wear", which does not directly relate to the target question, causing the model to erroneously infer the wrong label. While CoT successfully determines the correct answer, it fails to acknowledge the mention of "Amish cousins", thus exhibiting a tenuous connection to the question.
Unexpectedly, Self-Refine underperforms compared to other baselines, aligning with Huang et al. (2023) who highlight the drawbacks of self-correction in reasoning tasks. The primary challenge stems from the intricacy of designing few-shot examples that can effectively drive successive enhancements over prior outputs, limiting the potential for self-improvement. SEA-CoT, however, not only prompts self-assessment but also offers targeted guidance to enhance reasoning consistency with the relevant context. This simple extension greatly improves the quality of the explanation, with no downside on performance.
@@ 6.2 Ablation
Table 1: Ablation over various ways in selecting reasoning steps to serve as an explanation, implemented on StrategyQA. (O&E) is the proposed SEA-CoT which uses both components.
This ablation seeks to study the effectiveness of choosing the most suitable reasoning chain. We break down SEA-CoT’s ranking components and assess each of them, namely the entailment and keyword overlapping score. We additionally implement a baseline of SC-CoT that randomly picks from the list of explanations corresponding to the majority answer. The results from Table 1 demonstrate the efficacy of considering both components of SEA-CoT when ranking reasoning explanations. Choosing the most probable reasoning step has shown to not perform well, whereas our approach targeted at enhancing the important traits of an explanation is simple and yet does not hinder performance. We also conduct additional studies on different values of N sequences in Table 3.
@@ 6.3 Model size
The scaling laws of model size primarily concern the downstream performance of LLMs but little is known regarding the influence on interpretability properties. We replicate the experiments on the StrategyQA dataset with a focus on SEA-CoT prompting. We present the results in Table 2. The largest model, 70B generally outperforms the smaller sizes across all metrics while observing the same phenomenon in mistake insertion, previously discussed in 6.1. The improvement over smaller sizes may also be attributed to the enhanced accuracy in generating entailment scores for the explanation, analogous to observing greater performance of larger models in NLI tasks. Llama-13B surprisingly performs worse than its smaller variant, despite having a bigger network. More importantly, we note that by using SEA-CoT, even a 7B-sized model can generate more interpretable reasoning chains than a 70B model with other baseline prompts.
@@ 7 Related Works
Interpretable CoT: Since its introduction, CoT has garnered interest in the research community to innovate adaptation of it Chu et al. (2023). Despite CoT being primarily introduced to facilitate better reasoning skills out of LLMs, there is much interest to see if these reasoning steps could be used to explain the model’s thought process. Most of such works primarily investigate the faithfulness of the reasoning Lanham et al. (2023); Radhakrishnan et al. (2023); Turpin et al. (2023) or improving the faithfulness in CoT outputs, via refinement through knowledge retrieval He et al. (2022), symbolic reasoning Lyu et al. (2023), iterative information selection Creswell and Shanahan (2022) and factuality calibration Ye and Durrett (2022).
Concurrently, there are works Wang et al. (2023); He et al. (2022); Prasad et al. (2023) that focus on ascertaining the faithfulness of an explanation to the presence of factuality. While factuality is an important trait, it is not a sufficient component to ascertain faithfulness. Non-factual explanations may still align faithfully with an incorrect answer. Other works concentrate on semantic correctness Golovneva et al. (2022), regarded closer to plausibility, which differs from the traits assessed in this study. Our work strives to conduct a holistic assessment of interpretability across various forms of prompting techniques used in LLMs, taking into account multiple important properties that may be of importance to various audiences.
@@ 8 Conclusion
This work studied multiple ways of assessing the interpretability of an explanation. Our work is centered around assessing different variants of CoT and how we can better determine the suitability of the reasoning by-product as an explanation for the underlying prediction. We also propose a modification to the SC-CoT framework called SEA-CoT, designed specifically to yield explanations that better fulfill the objectives of interpretability. Our proposed framework surpasses the Robustness, Faithfulness, and Utility dimensions across multiple reasoning benchmarks. In the future, we plan to extend our work towards instilling interpretability and safety in the training stages Yang et al. (2023), such as safety alignment in LLM.
@@ 9 Limitations
Our work only investigates a single LLM - Llama-2 This work could be extended toward transformers of different structures such as encoder or encoder-decoder, or larger models, such as GPT3.5/4.0, which due to limiting resources are restricted to generate modifications instead. A secondary limitation is the quality of modifications to the original explanation, though we ask the modifier to check the outcome of the modified inputs (ie output remains the same when paraphrased), the correctness is nonetheless subjected to the ability of the modifier. Lastly, this work left out other techniques such as grounding the LLM’s response via external knowledge, which we note is an interesting avenue to consider next.
@@ A.2 Inference details
As we do not use API for the bulk of the experiments except perturbation generation and ablation using GPT3-5. We mainly rely on local resources to conduct inference. We use 4 x A6000 GPU for all experiments, each GPU has 46GB of VRAM and this gives us a total of 184GB VRAM. A 70B model would require at least 140GB VRAM, leaving only 44 VRAM left for text generation. Given an average input size of 1000 (usually longer for prompts such as QD) and a single batch size of 1, it would require an additional >60 GB VRAM (computed based on L = 80, H= 64, dim = 8192 for 70B) which makes it infeasible to implement. Thus, we perform the experiments using a 4-bit quantized version instead, which is performed using GPTQ on the original Llama-2 70B model. GPTQ is suitable for quantizing models consisting of billions of parameters. It has been validated on models up to 176B parameters and shown comparable performance with 16-bit models. The GPTQ-ed models are readily available on huggingface.
@@ A.4 Few-shot Prompts
For Self-Refine, there are three stages of instruction-prompting, where the second (feedback) and third (refine) stages continue iteratively until the LLM detects a stopping criterion which ends the cycle, denoted as "Stop refining the answer.". In the initial generation, the optimal examples are given, similar to CoT. In the feedback stage, we list scoring criteria which is focused on improving the interpretability of the reasoning explanation, instead of focusing on the performance. To simulate various qualities of output, we include both positive and negative examples. The examples in the refine stage are similar to the feedback but are instead designed in a continuous conversion displaying the full process of refining a bad example into a good one. We limit the number of examples in the refine stage to 3 as the context length is much longer here. The few-shot example prompts are displayed from Figure 12 to 15.
@@ A.7 Leakge-Adjusted Simulatability (LAS)
The first scenario can be solved by introducing a baseline, $\hat{y}|x$, and subtracting the task performance, $\mathbbm{1}[\hat{y}_{i}|\hat{e},x_{i}]$ from $\mathbbm{1}[\hat{y}_{i}|x_{i}]$. The second case is accounted by measuring the performance when the explanation either leaks, $\mathbbm{1}[\hat{y}_{i}|\hat{e}]=1$ or not, $\mathbbm{1}[\hat{y}_{i}|\hat{e}]=0$. $\mathbbm{1}$ denotes the event where outcome is correctly guessed, $\hat{y}=y$.

### v3

@@ Abstract
Prompt Engineering has garnered significant attention for enhancing the performance of large language models across a multitude of tasks. Techniques such as the Chain-of-Thought not only bolster task performance but also delineate a clear trajectory of reasoning steps, offering a tangible form of explanation for the audience. Prior works on interpretability assess the reasoning chains yielded by Chain-of-Thought solely along a singular axis, namely faithfulness. We present a comprehensive and multifaceted evaluation of interpretability, examining not only faithfulness but also robustness and utility across multiple commonsense reasoning benchmarks. Likewise, our investigation is not confined to a single prompting technique; it expansively covers a multitude of prevalent prompting techniques employed in large language models, thereby ensuring a wide-ranging and exhaustive evaluation. In addition, we introduce a simple interpretability alignment technique, termed Self-Entailment-Alignment Chain-of-thought, that yields more than 70% improvements across multiple dimensions of interpretability. Code is available at https://github.com/SenticNet/CoT_interpretability
@@
[…]
@@ 1 Introduction
In recent trends, Large Language Models (LLM) have shown impressive performance across a diverse array of tasks, primarily through extensive scaling of model size Brown et al. (2020). Techniques such as instruct-tuning Wei et al. (2021) applied across diverse tasks have empowered LLMs to execute inference on previously unseen tasks. One attributing factor lies with the extensive efforts put into innovating new ways of prompting the LLM to better exploit their knowledge base. Chain-of-Thought (CoT) Wei et al. (2022) has gathered much attention due to its simple setup which allows the LLM to generate not only the task output but also the steps undertaken.
In addition to its efficacy in enhancing the model’s performance, this prompting method concurrently touches on one of the important aspects of utilizing these models for decision-making: interpretability. The assumption is that the reasoning chain preceding the answer illustrates the model’s thought process, enabling the audience to understand how the answer is derived. However, such claims though seemingly plausible should be taken lightly as they may not be faithful to the model’s reasoning process Jacovi and Goldberg (2020). In this context, plausibility refers to the extent to which an explanation resonates with and is deemed acceptable by a human audience. Faithfulness, on the other hand, is characterized by the extent to which the explanation accurately reflects the model’s decision-making process.
There has been a large number of works that seek to introduce modifications to CoT, including Self-Consistency Wang et al. (2022b), Least-to-Most Zhou et al. (2022), while others specifically focus on establishing faithful reasoning Creswell and Shanahan (2022); Lyu et al. (2023). We introduce a simple extension to the list of CoT variants, but purely with a focus on enhancing interpretability in the reasoning chain. The approach coined Self-Entailment-Alignment CoT (SEA-CoT) operates similarly to Self-Consistency, but additionally utilizes a form of consistency between the corresponding reasoning steps and supported context. This action is missing in Self-Consistency, as the focus is only on the resultant output, potentially leading to unfaithful reasoning which may not support the underlying answer.
Moreover, we conduct an extensive investigation into the reasoning explanations by evaluating under three pivotal axes of interpretability: faithfulness, robustness, and utility on three commonsense reasoning datasets. These assessments are implemented across multiple prompting techniques including CoT and various adaptations of it.
@@ 3 Prompt Techniques
Figure 2: Overview of different prompting techniques to derive the reasoning chain, to serve as the explanation (boxed with dashed line). [Left to Right]: Cot, SC-CoT, SEA-CoT, QD, and Self-Refine (SR). SC-CoT and SEC-CoT differ in the explanation selection stage, where the former selects based on maximum cumulative probability and the latter on two objectives: entailment, $E$, and overlap, $O$ with an additional forward pass. Each robot figure denotes a forward pass from the LLM, SR stops when a stopping criteria is encountered or exceeds the max number of passes. SR requires the most pass, 3 per round.
@@ 4 Proposed Approach
Most adaptations on CoT are only aimed at maximizing task performance as covered in Section 3. Our work is instead focused on enhancing the interpretability of the presented reasoning chain preceding the task output. We adapt from SC-CoT, by focusing on the $N$ sequences produced, ranked based on specified objectives. Instead of picking explanations based on heuristics such as highest cumulative probability, the reasoning is chosen based on the maximization of two objectives: entailment and overlapping score between the supported context $(x\oplus\hat{y})$ and reasoning $\hat{e}$.
We posit that a credible explanation should intrinsically align with the given context it aims to elucidate Jie et al. (2024); in this scenario, it encompasses both the question being addressed and the prediction label, measured by the level of entailment.
We additionally maximize the overlap between two sets of key tokens [^1] The two sets of tokens are compared after removing any stopwords to minimize noise within the context, which we show in later experiments to be beneficial towards producing higher quality explanations. This simple approach can be regarded as performing a self-alignment step to pick the most suitable explanation with the $N$ sequences.
@@ 5 Interpretability Qualities
Utility: A largely understudied but important trait, utility is paramount to maximizing the information conveyed to the audience. A useful explanation can allow the discovery of new knowledge to human users such as understanding the causal relationships or enable more efficient knowledge distillation between neural models.
@@ 5.4 Counterfactual reasoning
An alternative method to ascertain faithfulness follows by evaluating whether an explanation would change when the original question is modified in a different direction, particularly when directed towards a counterfactual scenario. Atanasova et al. (2023) shows that an instance of unfaithfulness can be detected if the counterfactual explanation, $e^{\prime}$ does not acknowledge the modifications, $c$ in the counterfactual instance $x_{i}^{\prime}:y^{\prime}$, yet still successfully predicting the counterfactual label, $y^{\prime}\neq y$. The distinction from Section 5.2 is that besides detecting signs of unfaithfulness, it also embodies a directed approach that assesses a model’s capacity to contemplate alternative scenarios.
Conversely, introducing mistakes can be seen as an undirected measure aimed at gauging the decline in confidence, given an erroneous prior belief. We deemed an instance of unfaithfulness under the following conditions:
@@ 6 Experiments
Datasets: We implement perturbation experiments across three commonsense reasoning benchmarks.
@@ 6.1 Results
The key distinction between SC-CoT and SEA-CoT is the latter’s self-critique step, which evaluates how its explanations align with the context and the intended answer. This approach significantly boosts utility and reduces unfaithfulness in counterfactual contexts. Higher utility scores support the hypothesis that context-aligned stimuli enhance the efficiency of learning signals, facilitating easier training for student models.
Looking closer in Figure 6, where the word "shunned" is mentioned while other baselines used "would not wear", which does not directly relate to the target question, causing the model to erroneously infer the wrong label. While CoT successfully determines the correct answer, it fails to acknowledge the mention of "Amish cousins", thus exhibiting a tenuous connection to the question. Unexpectedly, Self-Refine underperforms compared to other baselines, aligning with Huang et al. (2023) who highlight the drawbacks of self-correction in reasoning tasks.
The primary challenge stems from the intricacy of designing few-shot examples that can effectively drive successive enhancements over prior outputs, limiting the potential for self-improvement. SEA-CoT, however, not only prompts self-assessment but also offers targeted guidance to enhance reasoning consistency with the relevant context. This simple extension greatly improves the quality of the explanation, with no downside on performance.
@@ 6.2 Ablation
Table 1: Ablation over ways of selecting reasoning steps to serve as an explanation on StrategyQA. (O&E) is the proposed SEA-CoT which uses both components.
This ablation seeks to study the effectiveness of choosing the most suitable reasoning chain. We break down SEA-CoT’s ranking components and assess each of them, namely the entailment and keyword overlapping score. We additionally implement a baseline of SC-CoT that randomly picks from the list of explanations corresponding to the majority answer. The results from Table 1 demonstrate the efficacy of considering both components of SEA-CoT when ranking reasoning explanations.
Choosing the most probable reasoning step has shown to not perform well, whereas our approach targeted at enhancing the important traits of an explanation is simple and yet does not hinder performance. We also conduct additional studies on different values of N sequences in Table 3.
@@ 6.3 Model size
The scaling laws of model size primarily concern the downstream performance of LLMs but little is known regarding the influence on interpretability properties. We replicate the experiments on the StrategyQA dataset with a focus on SEA-CoT prompting. We present the results in Table 2. The largest model, 70B generally outperforms the smaller sizes across all metrics while observing the same phenomenon in mistake insertion, previously discussed in 6.1. The improvement over smaller sizes may also be attributed to the enhanced accuracy in generating entailment scores for the explanation, analogous to observing greater performance of larger models in NLI tasks. Llama-13B surprisingly performs worse than its smaller variant, despite having a bigger network. Moreover, we note that by using SEA-CoT, even a 7B-sized model can generate more interpretable reasoning chains than a 70B model with other baseline prompts.
@@ 7 Related Works
Interpretable CoT: Since its introduction, CoT has garnered interest in the research community to innovate adaptation of it Chu et al. (2023). Despite CoT being primarily introduced to improve reasoning skills of LLMs, there is much interest to see if these reasoning steps could be used to explain the model’s thought process. Most works primarily investigate faithfulness of the reasoning Lanham et al. (2023); Radhakrishnan et al. (2023); Turpin et al. (2023) or improving the faithfulness in CoT outputs, via refinement through knowledge retrieval He et al. (2022), symbolic reasoning Lyu et al. (2023), iterative information selection Creswell and Shanahan (2022) and factuality calibration Ye and Durrett (2022). Other works focus on ascertaining the faithfulness of an explanation to the presence of factuality Wang et al. (2023); He et al. (2022); Prasad et al. (2023). While factuality is an important trait, it is not a sufficient component to ascertain faithfulness. Non-factual explanations may still align faithfully with an incorrect answer. Other works concentrate on semantic correctness Golovneva et al. (2022), regarded closer to plausibility, which differs from the traits assessed in this study. Our work strives to conduct a holistic assessment of interpretability across various forms of prompting techniques used in LLMs, taking into account multiple properties that may be of importance to various audiences.
@@ 8 Conclusion
This work studied multiple ways of assessing the interpretability of an explanation. Our work is centered on assessing different variants of CoT and how we can better determine the suitability of the reasoning by-product as an explanation for the underlying prediction. We also propose a modification to the SC-CoT framework called SEA-CoT, designed specifically to yield explanations that better fulfill the objectives of interpretability. Our proposed framework surpasses the Robustness, Faithfulness, and Utility dimensions across multiple reasoning benchmarks. In the future, we plan to extend our work towards instilling interpretability and safety in the training stages Yang et al. (2023), such as safety alignment in LLM.
@@ 9 Limitations
Our work only investigates a single LLM - Llama-2 This work could be extended toward transformers of different structures such as encoder or encoder-decoder, or larger models, such as GPT3.5/4.0, which due to limiting resources are restricted to generate modifications instead. A secondary limitation is the quality of modifications to the original explanation, though we ask the modifier to check the outcome of the modified inputs (i.e. output remains the same when paraphrased), the correctness is nonetheless subjected to the ability of the modifier. This work did not study techniques that ground the LLM’s response via external knowledge, which we note is an interesting avenue to consider next. An inherent weakness in LLM is self hallucination where it produces plausible text which are non-factual. Our work also left out investigating hybrid approaches such as Neuro-symbolic AI, which combines the learning abilities of neural networks and inherent interpretable decision-making frameworks of symbolic AI.
@@ A.2 Inference details
We do not use API for the bulk of the experiments except perturbation generation and ablation using GPT3-5. We mainly rely on local resources to conduct inference. We use 4 x A6000 GPU for all experiments, each GPU has 46GB of VRAM and this gives us a total of 184GB VRAM. A 70B model would require at least 140GB VRAM, leaving only 44 VRAM left for text generation. Given an average input size of 1000 (usually longer for prompts such as QD) and a single batch size of 1, it would require an additional >60 GB VRAM (computed based on L = 80, H= 64, dim = 8192 for 70B) which makes it infeasible to implement. Thus, we perform the experiments using a 4-bit quantized version instead, which is performed using GPTQ on the original Llama-2 70B model. GPTQ is suitable for quantizing models consisting of billions of parameters. It has been validated on models up to 176B parameters and shown comparable performance with 16-bit models. The GPTQ-ed models are readily available on huggingface.
@@ A.4 Few-shot Prompts
For Self-Refine, there are three stages of instruction-prompting, where the second (feedback) and third (refine) stages continue iteratively until the LLM detects a stopping criterion which ends the cycle, denoted as "Stop refining the answer.". In the initial generation, the optimal examples are given, similar to CoT. In the feedback stage, we list scoring criteria which are focused on improving the interpretability of the reasoning explanation, instead of focusing on the performance. To simulate various qualities of output, we include both positive and negative examples. The examples in the refine stage are similar to the feedback but are instead designed in a continuous conversion displaying the full process of refining a bad example into a good one. We limit the number of examples in the refine stage to 3 as the context length is much longer here. The few-shot example prompts are displayed from Figure 12 to 15.
@@ A.7 Leakge-Adjusted Simulatability (LAS)
The first scenario can be solved by introducing a baseline, $\hat{y}|x$, and subtracting the task performance, $\mathbbm{1}[\hat{y}_{i}|\hat{e},x_{i}]$ from $\mathbbm{1}[\hat{y}_{i}|x_{i}]$. The second case is accounted for by measuring the performance when the explanation either leaks, $\mathbbm{1}[\hat{y}_{i}|\hat{e}]=1$ or not, $\mathbbm{1}[\hat{y}_{i}|\hat{e}]=0$. $\mathbbm{1}$ denotes the event where outcome is correctly guessed, $\hat{y}=y$.